# On the Welfare of EIP-1559 with Patient Bidders

### v1

@@ Abstract
The “EIP-1599 algorithm” is used by the Ethereum blockchain to assemble transactions into blocks. While prior work has studied it under the assumption that bidders are “impatient”, we analyze it under the assumption that bidders are “patient”, which better corresponds to the fact that unscheduled transactions remain in the mempool and can be scheduled at a later time. We show that with “patient” bidders, this algorithm produces schedules of near-optimal welfare, provided it is given a mild resource augmentation (that does not increase with the time horizon). We prove some generalizations of the basic theorem, establish lower bounds that rule out several candidate improvements and extensions, and propose several questions for future work.
@@ 1.1 Background on Blockchains
Blockchains like Bitcoin [Nak09] or Ethereum [But14] operate by repeatedly aggregating transactions into blocks. As a first approximation, each transaction requires some “size” and has some value for its user, and each block has some size limit, where the notion of “size” is defined by the blockchain. [^3] E.g. “gas” for Ethereum or (virtual-)bytes for Bitcoin. Each block is assembled by a different “operator” [^4] In Bitcoin these are the “miners” and in Ethereum these are the “validators”. which the blockchain protocol designer tries to motivate to maximize the total value subject to the block’s capacity constraints. A key difficulty here is that both users and operators are selfish and strategic, so the mechanism that specifies which transactions are accepted to the block and how much do their users pay and to whom must take such strategic behavior into account. A large literature has studied such mechanisms. While Bitcoin’s “pay your bid” mechanisms is relatively simple (but not incentive compatible for the users), Ethereum’s mechanism [BCD+19], known as EIP-1559, is more sophisticated and, following [Rou21], a large literature (e.g. [LMR+21, CS23, GTW24, AGHH23]) studies its strategic properties.
@@
In this paper we undertake an algorithmic analysis of this protocol and the family of protocols that vary its various parameters. Specifically, we ask to what extent does it indeed succeed in maximizing the social welfare given the size constraints. Our main result is essentially positive: for “patient bidders”, the social welfare achieved by EIP-1559, with a slight tweak of parameters, is close to optimal. Before specifying more precisely what “close to optimal” means, we need to be more specific about our time-model of transactions.
@@
Blocks are created one at a time [^5] For Ethereum a block every 12 seconds and for Bitcoin a block every 10 minutes on the average. and transactions arrive continuously through time. Much analysis of blockchains assumes that a transaction that arrives at a certain point in time must be scheduled in the immediately next block or loses all value for its user. This is called the model of “impatient users”. In reality, however, blockchains operate under an opposite assumption and submitted transactions remain in the “mempool” until accepted, and thus may be scheduled at any future block. The model that assumes that transactions can be scheduled at any future time without losing their value, which is closer to how blockchains actually operate, is termed a model with “patient users”, and has only received modest attention [Nis23, HLM21, PS24, GY24]. A more general model may consider a mix of patient, impatient, and also “partially patient” users, who may have some time-sensitivity, e.g., model the value of a transaction as decreasing with its execution time [HLM21, PS24, GY24]. Unfortunately, current mainstream blockchain transaction fees mechanisms do not allow the user to express any time preferences. We show the near-optimally of EIP-1559 for fully patient users in the worst case, i.e. without any stochastic or distributional assumptions on the demand.666 It is not difficult to observe that if one looks, instead, at impatient users, then any algorithm that relies on historical prices to set inclusion criteria for the current block (as does EIP-1559) may catastrophically fail to achieve any decent social welfare under adversarial conditions. We will also show that no online algorithm can achieve near-optimally for a mix of patient and either impatient or partially patient users (even if they are “very patient”); see Section 1.5.1.
@@ 1.2 The EIP-1559 algorithm and its parameters
The EIP-1559 algorithm assembles a sequence of blocks whose average size is at a target of a quantity $B$ per block. [^7] Ethereum’s choice is $B=15M$ gas. I.e. the original constraint of maximum block size is relaxed, and is replaced by effectively limiting the average block size to be $B$, while relaxing the constraint of maximum block size to being only slightly larger than $B$ [^8] Maximum size of only $2B$, a factor $2$ increase that is independent of the parameters of the input (such as the range of per-unit values and the time horizon). This relaxation is justified by the specific constraints of the Ethereum blockchain [^9] I.e. it was deemed that the “syncing time” of “full nodes” (i.e., reading the entire history of the block chain) is more significant than the maximum block size limitation, and for it, only the average block size is important. and may be viewed as a type of resource augmentation which was not technically studied in the literature of bin-packing and knapsack algorithms, and may be of independent interest. [^10] This relaxation may be related in motivation to open-ended bin packing (e.g. [Eps22]), but not technically so.
[…]
@@
This algorithm produces blocks whose size may vary in the range 00 to $c\cdot B$. The target block size $B$ is reached only in an average sense, but in a rather strong sense of averaging that we wish to formalize as follows.
@@
A schedule is said to have average block size limit $B$ with slackness $\Delta:\mathbb{N}\rightarrow\mathbb{N}$ if for any number of times steps $k$, the total size of transactions during any $k$ consecutive blocks is at most $(k+\Delta(k))\cdot B$. When $\Delta$ is a constant function (for constant $Z$ it holds that $\Delta(k)=Z$), we abuse notation and simply write $\Delta=Z$.
The basic property of the EIP-1559 algorithm is that it indeed produces blocks whose average size is the target size $B$ [^17] Our analysis is for the theoretically clean version with $p_{t+1}=p_{t}\cdot e^{(\eta(Q_{t}-B)/B)}$. As shown in [LRMP23], the variant with $p_{t+1}=p_{t}\cdot(1+\eta(Q_{t}-B)/B)$ results in a slightly larger average block size. While this general fact is well known, we formalize it in our notion of “average block size limit $B$ with slackness $\Delta$” as follows:
@@ 1.3 Our Main result
Our main result states that the EIP-1599 algorithm is near-optimal for patient bidders, under some mild relaxations. Specifically, its allocation over slightly larger time horizon has social welfare that is close to that of the optimal, clairvoyantly chosen, schedule, that has maximum block size $B$.
@@
Such algorithms must “search” for the correct prices, which may be difficult under adversarial conditions. As we show, the EIP-1599 algorithm manages to do so with a loss that can be overcome by extending the time horizon by the number of blocks that is logarithmic in the per-unit value range. This single logarithmic loss is global over the whole range of blocks, even though prices may keep fluctuating.
@@
We thus see that the EIP-1559 algorithm can indeed compete with the optimal schedule, but after suffering two types of losses: first, there is a loss of a small $\delta\leq\eta$ fraction of the social welfare, where as the adjustment rate $\eta$ approaches 00 this loss vanishes. Second, the algorithm is given $\Gamma$ additional rounds, where $\Gamma$ is inversely proportional to $\eta$. Conceptually, if $T$ was given in advance we could balance between these two losses (for any fixed $v_{max},p_{min}$ and $c$) by choosing $\eta=O(1/\sqrt{T})$, which would lose $O(1/\sqrt{T})$ fraction of welfare every step as well as require an additional $O(1/\sqrt{T})$ fraction of steps.181818 This is somewhat in the vein of regret bounds that balance between losing a small factor (that is proportional to the learning rate) every step and suffering an additional additive loss that is inversely proportional to the learning rate. Unlike in classic regret bounds the theorem does not combine both losses into a single $O(1/\sqrt{T})$ regret measure.
@@
    Maximum Block Size vs. Maximum Transaction Size: The limitation on the maximum block size $c>2$ is required only when the largest transaction size $q_{max}$ can be as large as the full target block size $B$. If we have a better bound on the maximum transaction size, then the limitation on $c$ can be relaxed to $c>1+q_{max}/B$, and in the bound on $\Gamma$ the expression $1/(c-2)$ is replaced by $1/(c-1-q_{max}/B)$ (with a slight change to the additive term $c$ as well). Notice that when we have any bound on the maximum transaction size $q_{max}/B<1$ then the actual choice made in Ethereum of $c=2$ suffices for obtaining a good approximation and losing only $\eta=1/8$ fraction of social welfare when allowed a small number of extra blocks.
@@ 1.4 Lower Bounds
Let us now look deeper at the losses as stated in the main theorem. Not only will we show that EIP-1559 indeed may suffer such losses, but we will also prove that such losses must be suffered by two classes of algorithms of which EIP-1599 is a member: a loss for the general class of online algorithms and an additional loss for a subclass of online algorithms that we term “price-based algorithms” that operate by setting a price for each block that only depends on on-chain information from previous blocks (see Definition 2.9 for a formal definition).
@@ 1.4.1 Maximum Block size
Any online algorithm that produces schedules with maximum block size $c\cdot B$, for some $c<2$ and average block size limit $B$ with slackness $\Delta(T)=o(T)$, even with an extension $\Gamma(T)=o(T)$, must lose at least $\delta_{c}=min(1/8,(2-c)/3)$ fraction of social welfare relative to schedules with worst-case block size limit $B$ (even when the values of all transactions are in $\{1,2\}$).
@@ 1.4.2 Dependence on Range of per-unit Values
Note that there is an exponential gap between our lower bound (which grows at rate $\Omega(\log\log H)$) and the upper bound (that is only singly-logarithmic in the per-unit value range). While we do show that the EIP-1559 algorithm may indeed require such a singly-logarithmic additional number of blocks in order to compete with the optimum, closing the gap for general price-based algorithms remains open. Especially intriguing is the possibility of designing a useful “variant” of EIP-1559 with $\Delta+\Gamma$ that only grows in a rate double-logarithmic in the per-unit value range.
@@ 1.5.1 Partially Patient Bidders
All our results so far assumed “patient bidders”, i.e. where a transaction’s value for its user remains the same over time. The opposite assumption of “impatient bidders” assumes that a transaction must be scheduled in the immediate block or it loses all value for its user. A more general model would capture some sensitivity of the value of a transaction to the time of its execution, where the value decreases with time. The most common model for such time-dependent value [HLM21, PS24, GY24] would have a discount rate: A transaction with value $v$ and discount factor $\rho$ (where $10\leq\rho<$ 1) has value for its user of $v\cdot(1-\rho)^{t_{e}-t_{a}}$, where $t_{a}$ is its arrival time and $t_{e}$ its execution time.
@@
Fix an online algorithm with average block size $B$, slackness $\Delta$, and extension $\Gamma$, where $\Delta(T)+\Gamma(T)=o(T)$. For every minimum discount rate $\rho_{min}>0$ there exist a time horizon $T$, and an input sequence where every transaction $i$ has a discount rate $\rho_{i}\in\{0,\rho_{min}\}$ and a value in $v_{i}\in\{1,2\}$, for which the algorithm loses at least a fraction $\delta_{0}=1/20$ of welfare relative to the optimal schedule with worst-case block size limit $B$.
@@ 1.5.2 Multi-dimensional Fees
If there are at least three resources then any online algorithm with average block size $B_{j}$, slackness $\Delta(T)=o(T)$ for each resource $j$ and given an extension $\Gamma(T)=o(T)$, must lose at least a constant fraction $\delta_{0}=1/6$ of social welfare on some input sequence.
@@ 1.6 Strategic Points of View
Our analysis in this paper is purely algorithmic, so our results can be viewed as proving near-optimally in terms of the declared social welfare: the welfare with respect to the values given as input by the bidders to the algorithm (rather than the true values). Regarding the operators (miners, validators), our analysis assumed adversarial behavior, subject to the weak constraint that they must schedule a maximal by inclusion subset of admissible transactions. While we do not formally make any strategic analysis, we do want to shortly mention some strategic points of view.
From the point of view of the bidders, the known incentive compatibility of EIP-1559 is only for impatient bidders [^20] Even for impatient bidders, the incentive compatibility of EIP-1559 does not hold in the case of a rapid demand increase as then tips kick in and winners pay the tips they offer., while patient bidders may in fact shade their bids profitably. So can we expect truthful bidding from patient bidders? Well, practically, bidders may simply act myopically – in our case, truthfully – due to the complexity of predicting the future, and their bounded rationality. In our case, such behavior is also justified by risk aversion: Bidders that are sufficiently risk-averse with respect to the future will bid truthfully, taking the first available slot at a profit rather than risk never being scheduled due to possible future increase in demand. An appealing strategic analysis of patient bidders’ behavior that includes a reasonable model of future uncertainty is of course of much interest, and is left as a future challenge.212121 It is interesting to note that if in equilibrium the declared values are monotone in the true values – as may be expected in many models – then optimizing for the former is closely related to optimizing for the latter. In that case, our algorithmic results may carry over to the strategic setting in the model in question. A precise statement to such a result would obviously depends on many details of the model.

### v2

@@ Abstract
The “EIP-1559 algorithm” is used by the Ethereum blockchain to assemble transactions into blocks. While prior work has studied it under the assumption that bidders are “impatient”, we analyze it under the assumption that bidders are “patient”, which better corresponds to the fact that unscheduled transactions remain in the mempool and can be scheduled at a later time. We show that with “patient” bidders, this algorithm produces schedules of near-optimal welfare, provided it is given a mild resource augmentation (that does not increase with the time horizon). We prove some generalizations of the basic theorem, establish lower bounds that rule out several candidate improvements and extensions, and propose several questions for future work.
@@ 1.1 Background on Blockchains
Blockchains like Bitcoin [Nak09] or Ethereum [But14] operate by repeatedly aggregating transactions into blocks. As a first approximation, each transaction requires some “size” and has some value for its user, and each block has some size limit, where the notion of “size” is defined by the blockchain. [^3] E.g. “gas” for Ethereum or (virtual-)bytes for Bitcoin. Each block is assembled by a different “operator” [^4] In Bitcoin these are the “miners” and in Ethereum these are the “validators”. which the blockchain protocol designer tries to motivate to maximize the total value subject to the block’s capacity constraints. A key difficulty here is that both users and operators are selfish and strategic, so the mechanism that specifies which transactions are accepted to the block and how much do their users pay and to whom must take such strategic behavior into account. A large literature has studied such mechanisms. While Bitcoin’s “pay your bid” mechanism is relatively simple (but not incentive compatible for the users), Ethereum’s mechanism [BCD+19], known as EIP-1559, is more sophisticated and, following [Rou21], a large literature (e.g. [LMR+21, CS23, GTW24, AGHH23]) studies its strategic properties.
@@
In this paper we undertake an algorithmic analysis of this protocol and the family of protocols that vary its various parameters. Specifically, we ask to what extent it succeeds in maximizing the social welfare given the size constraints. Our main result is essentially positive: for “patient bidders”, the social welfare achieved by EIP-1559, with a slight tweak of parameters, is close to optimal. Before specifying more precisely what “close to optimal” means, we need to be more specific about our time-model of transactions.
@@
Blocks are created one at a time [^5] For Ethereum a block every 12 seconds and for Bitcoin a block every 10 minutes on the average. and transactions arrive continuously through time. Much analysis of blockchains assumes that a transaction that arrives at a certain point in time must be scheduled in the immediately next block or loses all value for its user. This is called the model of “impatient users”. In reality, however, blockchains operate under an opposite assumption and submitted transactions remain in the “mempool” until accepted, and thus may be scheduled at any future block. The model that assumes that transactions can be scheduled at any future time without losing their value, which is closer to how blockchains actually operate, is termed a model with “patient users”, and has only received modest attention [Nis23, HLM21, PS24, GY24]. A more general model may consider a mix of patient, impatient, and also “partially patient” users, who may have some time-sensitivity, e.g., model the value of a transaction as decreasing with its execution time [HLM21, PS24, GY24]. Unfortunately, current mainstream blockchain transaction-fee mechanisms do not allow the user to express any time preferences. We show the near-optimality of EIP-1559 for fully patient users in the worst case, i.e., without any stochastic or distributional assumptions on the demand. [^6] It is not difficult to observe that if one looks, instead, at impatient users, then any algorithm that relies on historical prices to set inclusion criteria for the current block (as does EIP-1559) may catastrophically fail to achieve any decent social welfare under adversarial conditions. We will also show that no online algorithm can achieve near-optimality for a mix of patient and either impatient or partially patient users (even if they are “very patient”); see Section 1.5.1.
@@ 1.2 The EIP-1559 algorithm and its parameters
The EIP-1559 algorithm assembles a sequence of blocks with target average size of $B$ per block. [^7] Ethereum’s choice is $B=18M$ gas. I.e. the original constraint of maximum block size is relaxed, and is replaced by effectively limiting the average block size to be $B$, while relaxing the constraint of maximum block size to being only slightly larger than BB. [^8] Maximum size of only $2B$, a factor $2$ increase that is independent of the parameters of the input (such as the range of per-unit values and the time horizon). This relaxation is justified by the specific constraints of the Ethereum blockchain [^9] I.e. it was deemed that the “syncing time” of “full nodes” (i.e., reading the entire history of the block chain) is more significant than the maximum block size limitation, and for it, only the average block size is important. and may be viewed as a type of resource augmentation which was not technically studied in the literature of bin-packing and knapsack algorithms, and may be of independent interest. [^10] This relaxation may be related in motivation to open-ended bin packing (e.g. [Eps22]), but not technically so.
Here is how it operates: for every block $t$ it chooses a price $p_{t}$ (according to an algorithm that we will explain shortly) and accepts only transactions whose value per unit is at least the price $p_{t}$. A difficulty arises if there are too many transactions with such high value per unit, as the EIP-1559 algorithm does not allow any block to be of total size that is larger than $c$ times the target block size, with the parameter choice c=2c=2.111111I.e. the maximum allowed block size in Ethereum is $36M$ gas. In this case only a subset of the transactions that fit into the maximum block size is chosen, with the subset determined by an additional “tip” bid that is strategically chosen by each user. In this paper we will not undertake a strategic analysis, and we will analyze the protocol under the assumption that the subset of transactions is being chosen by an adversary [^12] This naturally only strengthens our main result., making only the following mild assumption: the adversary must pick a maximal (by inclusion) set of transactions that fit into the allowed maximum block size. [^13] This is a mild assumption as any rational operator that is myopic (cares only about the current block and disregards the future) is never better off by excluding transactions that can simply be added to the block. In particular, the presence of (non-negative) tips only encourages the inclusion of additional transactions in the block.
@@
This algorithm produces blocks whose size may vary in the range 0 to $c\cdot B$. The target block size $B$ is reached only in an average sense, but in a rather strong sense of averaging that we wish to formalize as follows.
@@
For a monotone non-decreasing function $\Delta:\mathbb{N}\rightarrow\mathbb{N}$, a schedule is said to have average block size limit $B$ with slackness $\Delta$, if for any number of times steps $k$, the total size of transactions during any $k$ consecutive blocks is at most $(k+\Delta(k))\cdot B$. When $\Delta$ is a constant function (for constant $Z$ it holds that $\Delta(k)=Z$), we abuse notation and simply write $\Delta=Z$.
The basic property of the EIP-1559 algorithm is that it indeed produces blocks whose average size is the target size BB. [^17] Our analysis is for the theoretically clean version with $p_{t+1}=p_{t}\cdot e^{(\eta(Q_{t}-B)/B)}$. As shown in [LRMP23], the variant with $p_{t+1}=p_{t}\cdot(1+\eta(Q_{t}-B)/B)$ results in a slightly larger average block size. While this general fact is well known, we formalize it in our notion of “average block size limit $B$ with slackness $\Delta$” as follows:
@@ 1.3 Our Main result
Our main result states that the EIP-1559 algorithm is near-optimal for patient bidders, under some mild relaxations. Specifically, its allocation over a slightly larger time horizon has social welfare that is close to that of the optimal, clairvoyantly chosen, schedule, that has maximum block size $B$.
@@
Such algorithms must “search” for the correct prices, which may be difficult under adversarial conditions. As we show, the EIP-1559 algorithm manages to do so with a loss that can be overcome by extending the time horizon by the number of blocks that is logarithmic in the per-unit value range. This single logarithmic loss is global over the whole range of blocks, even though prices may keep fluctuating.
@@
We thus see that the EIP-1559 algorithm can indeed compete with the optimal schedule, but after suffering two types of losses: first, there is a loss of a small $\delta\leq\eta$ fraction of the social welfare, where as the adjustment rate $\eta$ approaches 0 this loss vanishes. Second, the algorithm is given $\Gamma$ additional rounds, where $\Gamma$ is inversely proportional to $\eta$. Conceptually, if $T$ was given in advance we could balance between these two losses (for any fixed $v_{max},p_{min}$ and $c$) by choosing $\eta=O(1/\sqrt{T})$, which would lose $O(1/\sqrt{T})$ fraction of welfare every step as well as require an additional $O(1/\sqrt{T})$ fraction of steps. [^18] This is somewhat in the vein of regret bounds that balance between losing a small factor (that is proportional to the learning rate) every step and suffering an additional additive loss that is inversely proportional to the learning rate. Unlike in classic regret bounds the theorem does not combine both losses into a single $O(1/\sqrt{T})$ regret measure.
@@
    Maximum Block Size vs. Maximum Transaction Size: The limitation on the maximum block size $c>2$ is required only when the largest transaction size $q_{max}$ can be as large as the full target block size $B$. If we have a better bound on the maximum transaction size, then the limitation on $c$ can be relaxed to $c>1+q_{max}/B$, and in the bound on $\Gamma$ the expression 1/(c2)1/(c-2) is replaced by $1/(c-1-q_{max}/B)$ (with a slight change to the additive term $c$ as well). Notice that when we have any bound on the maximum transaction size $q_{max}/B<1$ then the actual choice made in Ethereum of $c=2$ suffices for obtaining a good approximation and losing only $\eta=1/8$ fraction of social welfare when allowed a small number of extra blocks.
@@ 1.4 Lower Bounds
Let us now look deeper at the losses as stated in the main theorem. Not only will we show that EIP-1559 indeed may suffer such losses, but we will also prove that such losses must be suffered by two classes of algorithms of which EIP-1559 is a member: a loss for the general class of online algorithms and an additional loss for a subclass of online algorithms that we term “price-based algorithms” that operate by setting a price for each block that only depends on on-chain information from previous blocks (see Definition 2.9 for a formal definition).
@@ 1.4.1 Maximum Block size
Any online algorithm that produces schedules with maximum block size $c\cdot B$, for some $c<2$ (with $c\geq 1$) and average block size limit $B$ with slackness $\Delta(T)=o(T)$, even with an extension $\Gamma(T)=o(T)$, must lose at least $\delta_{0}=\Omega(2-c)$ fraction of social welfare relative to schedules with worst-case block size limit $B$ (even when the values of all transactions are in $\{1,2\}$).
@@ 1.4.2 Dependence on Range of per-unit Values
Note that there is an exponential gap between our lower bound (which grows at rate $\Omega(\log\log H)$) and the upper bound (that is only singly-logarithmic in the per-unit value range). While we do show that the EIP-1559 algorithm may indeed require such a singly-logarithmic additional number of blocks in order to compete with the optimum, closing the gap for general price-based algorithms remains open. Especially intriguing is the possibility of designing a useful “variant” of EIP-1559 with $\Delta+\Gamma$ that only grows at a rate double-logarithmic in the per-unit value range.
@@ 1.5.1 Partially Patient Bidders
All our results so far assumed “patient bidders”, i.e. where a transaction’s value for its user remains the same over time. The opposite assumption of “impatient bidders” assumes that a transaction must be scheduled in the immediate block or it loses all value for its user. A more general model would capture some sensitivity of the value of a transaction to the time of its execution, where the value decreases with time. The most common model for such time-dependent value [HLM21, PS24, GY24] would have a discount rate: A transaction with value $v$ and discount factor $\rho$ (where $0\leq\rho<1$) has value for its user of $v\cdot(1-\rho)^{t_{e}-t_{a}}$, where $t_{a}$ is its arrival time and $t_{e}$ its execution time.
@@
Fix an online algorithm with average block size $B$, slackness $\Delta$, and extension $\Gamma$, where $\Delta(T)+\Gamma(T)=o(T)$. For every maximum discount rate ${\rho_{max}}>0$ there exist a time horizon $T$, and an input sequence where every transaction $i$ has a discount rate $\rho_{i}\in\{0,{\rho_{max}}\}$ and a value in $v_{i}\in\{1,2\}$, for which the algorithm loses at least a fraction $\delta_{0}=1/20-o(1)$ of welfare relative to the optimal schedule with worst-case block size limit $B$.
@@ 1.5.2 Multi-dimensional Fees
If there are at least three resources then any online algorithm with average block size $B_{j}$, slackness $\Delta(T)=o(T)$ for each resource $j$ and given an extension $\Gamma(T)=o(T)$, must lose at least a constant fraction $\delta_{0}=1/6-o(1)$ of social welfare on some input sequence.
After the initial version of this paper was published, [BEN25] extended the lower bound also for the case of two resources, and additionally, gave a constant factor approximation for the general case.
@@ 1.6 Strategic Points of View
Our analysis in this paper is purely algorithmic, so our results can be viewed as proving near-optimality in terms of the declared social welfare: the welfare with respect to the values given as input by the bidders to the algorithm (rather than the true values). Regarding the operators (miners, validators), our analysis assumed adversarial behavior, subject to the weak constraint that they must schedule a maximal by inclusion subset of admissible transactions. While we do not formally make any strategic analysis, we do want to shortly mention some strategic points of view.
From the point of view of the bidders, the known incentive compatibility of EIP-1559 is only for impatient bidders [^20] Even for impatient bidders, the incentive compatibility of EIP-1559 does not hold in the case of a rapid demand increase as then tips kick in and winners pay the tips they offer., while patient bidders may in fact shade their bids profitably. So can we expect truthful bidding from patient bidders? Well, practically, bidders may simply act myopically – in our case, truthfully – due to the complexity of predicting the future, and their bounded rationality. In our case, such behavior is also justified by risk aversion: Bidders that are sufficiently risk-averse with respect to the future will bid truthfully, taking the first available slot at a profit rather than risk never being scheduled due to possible future increase in demand. An appealing strategic analysis of patient bidders’ behavior that includes a reasonable model of future uncertainty is of course of much interest, and is left as a future challenge. [^21] It is interesting to note that if in equilibrium the declared values are monotone in the true values – as may be expected in many models – then optimizing for the former is closely related to optimizing for the latter. In that case, our algorithmic results may carry over to the strategic setting in the model in question. A precise statement to such a result would obviously depend on many details of the model.